find the value of x by solving the equation. show your work
(10+2x)(5+2x)= 126

Answer by Boreal(15235)

(10+2x)(5+2x)= 126
50+20x+10x+4x^2=126
4x^2+30x-76=0
divide by 2
2x^2+15x-38=0
(2x+19)(x-2)=0
x=-9.5, x=2
Try 2
14*9=126
Try -9.5
(-9)(-14)=126
solutions are 2, -9.5